# FedGCS: A Generative Framework for Efficient Client Selection in Federated Learning via Gradient-based Optimization

### v1

@@ 2.2 Generative Models
Recent advancements in LLMs and generative models mark a shift in the machine learning paradigm, attesting to the evolution from task-specific architectures Ning et al. (2022) to highly generalized generic models. These models, exemplified by the likes of GPT-3 Brown et al. (2020) in the language domain, along with DALL·E Ramesh et al. (2022) in the multi-modal domain, have driven the migration to generative models across various domains Gruver et al. (2023); Xiao et al. (2023); Wang et al. (2024). Generative models have showcased an unparalleled ability to encode multifaceted knowledge representations from expansive data into their parameters Brown et al. (2020). This encoded knowledge can be adeptly maneuvered to generate contextually coherent and semantically rich outputs, addressing a wide spectrum of downstream tasks Ning et al. (2021). Despite the wide range of applications in language and vision, we have not seen generative models applied extensively in decision-making tasks like the client selection task in FL.
@@ 3 Problem Formulation
In this section, we formulate the discrete client selection task in FL as a generative task and conduct gradient-based optimization in a continuous representation space to get the optimal selection. Firstly, at the beginning of a training round, given a candidate client device pool set $\mathcal{C}=\{c_{1},c_{2},\dots\,c_{J}\}$ of $J$ client devices, we use classic client selection algorithms to collect $n$ “selection-score” pair as training data for subsequent model training. We denote the collected records as =$\mathcal{R}=\{(\mathbf{s_{i}},p_{i})\}_{1}^{n}$, where $\mathbf{s_{i}}=\{s_{1},s_{2},\dots\,s_{T}\}$ is the client selection record made by a classic algorithm ($\mathbf{s_{i}}\subset\mathcal{C}$, $s_{1\leq i\leq T}\in\mathcal{C}$ is the selected device IDs, and $T$ is not a fixed number that is determined by the classical algorithm used for this selection), and $p_{i}$ is the corresponding comprehensive score of the FL performance for making the client selection $\mathbf{s_{i}}$. To characterize $p_{i}$ of a given $\mathbf{s_{i}}$, unlike previous FL approaches Li et al. (2019); Zhan et al. (2020); Wang et al. (2020) that focus on different metrics and typically consider only one or two metrics to guide the selection, in this paper, we simultaneously focus on: (1) good performance of the global model; (2) low processing latency; (3) low energy consumption to perform comprehensive and multi-dimensional optimization. Here, the comprehensive metric function $\mathit{M}$ of a client selection $\mathbf{s_{i}}$ is defined as:
@@ 4.1 Framework Overview
To ensure that subsequent models and algorithms perform well, we need to collect sufficient, diverse, comprehensive and high-quality “selection-score” pair data =$\mathcal{R}=\{(\mathbf{s_{i}},p_{i})\}_{1}^{n}$ as training data: training data is large enough to represent the entire distribution; training data includes high-performance selection cases as well as certain random exploration samples. Intuitively, we can use random selections and compute their corresponding comprehensive scores as training data. However, this strategy is inefficient because the number of random selections is exceptionally large and most of them are low-performance samples. Therefore, we also must consider the efficiency of our data collection process.
@@ 5.1 Experimental Setup
We run two heuristic-based methods—Oort and Favor, and one heuristic-based method—FedMarl 100 times each, totaling 300 runs, to collect “selection-score” pair data for subsequent model training. For data augmentation, each selection is randomly shuffled 25 times to model set order-independence. We adopt a single-layer LSTM as the Encoder and Decoder backbones, and two-layer feed-forward networks for the Evaluator. The hidden state sizes are 64 (Encoder), 64 (Decoder), and 200 (Evaluator), with a $32$-dimensional embedding for each device ID token. For FedGCS training, we set batch size$=1024$, learning rate$=0.001$, $\alpha=0.8$, and utilize the top-$25$ device selections as starting points for gradient optimization. In the FL setting, $J=100$ devices are available, with $T=10$ ($T=20$ for FedMarl) selected per round over $r=10$ rounds for GPT-2 ($r=50$ for others). Clients perform local $ep=5$ epoch training each round. The code is available at https://github.com/zhiyuan-ning/GenerativeFL.
@@ 5.2 Overall Performance
Schemes Acc. (%$\uparrow$ Latency(min)$\downarrow$ Energy(KJ)$\downarrow$

### v2

@@ 2.2 Generative Models
Recent advancements in LLMs and generative models mark a shift in the machine learning paradigm, attesting to the evolution from task-specific architectures Ning et al. (2022, 2025a, 2025b) to highly generalized generic models. These models, exemplified by the likes of GPT-3 Brown et al. (2020) in the language domain, along with DALL·E Ramesh et al. (2022) in the multi-modal domain, have driven the migration to generative models across various domains Gruver et al. (2023); Xiao et al. (2023); Wang et al. (2024). Generative models have showcased an unparalleled ability to encode multifaceted knowledge representations from expansive data into their parameters Brown et al. (2020). This encoded knowledge can be adeptly maneuvered to generate contextually coherent and semantically rich outputs, addressing a wide spectrum of downstream tasks Ning et al. (2021). Despite the wide range of applications in language and vision, we have not seen generative models applied extensively in decision-making tasks like the client selection task in FL.
@@ 3 Problem Formulation
In this section, we formulate the discrete client selection task in FL as a generative task and conduct gradient-based optimization in a continuous representation space to get the optimal selection. Firstly, at the beginning of a training round, given a candidate client device pool set $\mathcal{C}=\{c_{1},c_{2},\dots\,c_{J}\}$ of $J$ client devices, we use classic client selection algorithms to collect $n$ “selection-score” pair as training data for subsequent model training. We denote the collected records as $\mathcal{R}=\{(\mathbf{s_{i}},p_{i})\}_{1}^{n}$, where $\mathbf{s_{i}}=\{s_{1},s_{2},\dots\,s_{T}\}$ is the client selection record made by a classic algorithm ($\mathbf{s_{i}}\subset\mathcal{C}$, $s_{1\leq i\leq T}\in\mathcal{C}$ is the selected device IDs, and $T$ is not a fixed number that is determined by the classical algorithm used for this selection), and $p_{i}$ is the corresponding comprehensive score of the FL performance for making the client selection $\mathbf{s_{i}}$. To characterize $p_{i}$ of a given $\mathbf{s_{i}}$, unlike previous FL approaches Li et al. (2019); Zhan et al. (2020); Wang et al. (2020) that focus on different metrics and typically consider only one or two metrics to guide the selection, in this paper, we simultaneously focus on: (1) good performance of the global model; (2) low processing latency; (3) low energy consumption to perform comprehensive and multi-dimensional optimization. Here, the comprehensive metric function $\mathit{M}$ of a client selection $\mathbf{s_{i}}$ is defined as:
@@ 4.1 Framework Overview
To ensure that subsequent models and algorithms perform well, we need to collect sufficient, diverse, comprehensive and high-quality “selection-score” pair data $\mathcal{R}=\{(\mathbf{s_{i}},p_{i})\}_{1}^{n}$ as training data: training data is large enough to represent the entire distribution; training data includes high-performance selection cases as well as certain random exploration samples. Intuitively, we can use random selections and compute their corresponding comprehensive scores as training data. However, this strategy is inefficient because the number of random selections is exceptionally large and most of them are low-performance samples. Therefore, we also must consider the efficiency of our data collection process.
@@ 5.1 Experimental Setup
We run two heuristic-based methods—Oort and Favor, and one heuristic-based method—FedMarl 100 times each, totaling 300 runs, to collect “selection-score” pair data for subsequent model training. For data augmentation, each selection is randomly shuffled 25 times to model set order-independence. We adopt a single-layer LSTM as the Encoder and Decoder backbones, and two-layer feed-forward networks for the Evaluator. The hidden state sizes are 64 (Encoder), 64 (Decoder), and 200 (Evaluator), with a $32$-dimensional embedding for each device ID token. For FedGCS training, we set batch size=1024=1024, learning rate=0.001=0.001, $\alpha=0.8$, and utilize the top-$25$ device selections as starting points for gradient optimization. In the FL setting, $J=100$ devices are available, with $T=10$ ($T=20$ for FedMarl) selected per round over $r=10$ rounds for GPT-2 ($r=50$ for others). Clients perform local $ep=5$ epoch training each round. The code is available at https://github.com/zhiyuan-ning/GenerativeFL.
@@ 5.2 Overall Performance
[…]
@@ Acknowledgments
Appendix
Appendix A The Detailed Process of Data Collection
In our implementation, we use one heuristic-base selection method: Oort Lai et al. (2021), and two learning-based selection methods: Favor Wang et al. (2020) and FedMarl Zhang et al. (2022) as our training data collectors to obtain appropriate selections. They differ in the metrics they focus on and the specific manner of their realization.
For Oort, it optimizes device selection by defining the utility of client $i$ as the product of the statistical utility and the global system utility, which integrates training loss and latency to maximize per-unit-time efficiency.
$\operatorname{Util}(i)=\underbrace{\left|B_{i}\right|\sqrt{\frac{1}{\left|B_{i}\right|}\sum_{k\in B_{i}}\operatorname{Loss}(k)^{2}}}_{\text{Statistical utility }U(i)}\times\underbrace{\left(\frac{T}{t_{i}}\right)^{\textbf{1}\left(T<t_{i}\right)\times\alpha}}_{\text{Global sys utility }}$ (10)
For Favor, it employs a double DQN-trained DRL agent to select local weights based on accuracy, countering Non-IID challenges and boosting convergence, optimizing rewards represented by:
$R=\sum_{t=1}^{T}\gamma^{t-1}\left(\Xi^{\left(\omega_{t}-\Omega\right)}-1\right)$ (11)
For FedMarl, it employs multi-agent reinforcement learning for device selection in federated learning, optimizing accuracy, latency, and communication cost, with rewards per round defined as:
$r_{t}=w_{1}[U(Acc(t))-U(Acc(t-1))]-w_{2}H_{t}-w_{3}B_{t}$ (12)
We ran each of the three training data collectors 100 times, thus collecting 300 client selections. For those diverse, comprehensive, and high-quality selections, we devise the following formula that combines performance, latency, and energy consumption to obtain the corresponding comprehensive scores $p$ for each collected selection, and thus the final “selection-score” training data:
$p$ $=p_{Perf}\times(\frac{L}{p_{L}})^{\textbf{1}(L<p_{L})\times\alpha}\times(\frac{E}{p_{E}})^{\textbf{1}(E<p_{E})\times\beta}$ (13)
where:
$p_{L}$ $=\max_{0\leq i\leq T}\{[L_{comm,i}+L_{comp,i}*l_{ep}]\}$
$p_{E}$ $=\sum_{T}\{[E_{comm,i}+E_{comp,i}*l_{ep}]\}$
where $L$ and $E$ are the developer-preferred duration and energy budget of the devices, respectively. comm represents the communication process and comp represents the computation process. $l_{ep}$ is the number of local training epochs. $\textbf{1}(x)$ is an indicator function that takes the value 1 if $x$ is true and 0 otherwise. In this way, the utility of those clients that may be the bottleneck of the desired speed and energy cost of the current round is penalized by a developer-specified factor $\alpha$ and $\beta$, but we do not reward the non-straggler and non-overloader clients because their completions do not affect the effectiveness of the training round Lai et al. (2021). We set both $\alpha$ and $\beta$ to 2 in our implementation.
Appendix B Experiment Details
B.1 Dataset Details
We use four computer vision datasets and two natural language processing datasets to fully validate the capabilities of our methodology. Following are their detailed descriptions:
  • MNIST Lecun et al. (1998b) is a large collection of handwritten digits. It has a training set of $60,000$ examples and a test set of $10,000$ examples. The images are centered in a $28\times 28$ image by computing the center of mass of the pixels, and translating the image so as to position this point at the center of the $28\times 28$ field.
  • CIFAR10 Krizhevsky et al. (2009) is a subset of the Tiny Images dataset and consists of $60,000$ $32\times 32$ color images. These images are labeled into one of ten mutually exclusive categories: airplanes, automobiles (but not trucks or pickups), birds, cats, deer, dogs, frogs, horses, ships, and trucks (but not pickups). There are $6000$ images for each category, with 5000 training images and 1000 test images.
  • CINIC10 Darlow et al. (2018) is a dataset for image classification that has a total of $270,000$ images. It is constructed from two different sources: ImageNet and CIFAR10. It is split into three equal subsets—train, validation, and test—each of which contains $90,000$ images.
  • TinyImageNet Deng et al. (2009) contains $100,000$ images of $200$ classes ($500$ for each class) downsized to $64\times 64$ colored images. Each class has $500$ training images, $50$ validation images, and $50$ test images.
  • Shakespeare Shakespeare (2017) is a dataset built from The Complete Works of William Shakespeare. It is a popular choice for training language models due to its manageable size and the complexity of Shakespeare’s language. It provides a good balance between computational efficiency and the ability to generate interesting text.
  • WikiText Merity et al. (2016) is a language modeling dataset that is a collection of over 100 million tokens extracted from the set of verified Good and Featured articles on Wikipedia. The dataset is available under the Creative Commons Attribution-ShareAlike License.
For four computer vision datasets, we generate IID data splits by randomly assigning training examples to each client without replacement. For Non-IID splits, we simulate data heterogeneity by sampling label ratios from a Dirichlet distribution $p_{k}\sim Dir_{N}(\beta)$ with the symmetric parameter $\beta$ Hsu et al. (2019). We set $\beta=0.01$ for MNIST, and $\beta=0.1$ for others to emulate Non-IID. For two natural language processing datasets, Shakespeare and Wikitext are naturally non-IID. We use LeNet5 LeCun et al. (1998a) on MNIST, ResNet18 He et al. (2016) on CIFAR10, VGG16 Simonyan. (2014) on CINIC10, ShuffleNet Zhang et al. (2018) on TinyImageNet, LSTM Hochreiter and Schmidhuber (1997) on Shakespeare, GPT-2 Radford et al. (2019) on Wikitext. For a fair comparison, baselines and FedGCS will be compared under the same settings.
Dataset &Model CV Tasks NLP Tasks
MNIST CIFAR10 CINIC10 TinyImageNet Shakespeare Wikitext
LeNet5 Resnet18 VGG16 ShuffleNet LSTM GPT-2
FedAvg 9.84 10.21 28.24 31.67 67.89 71.23 81.43 85.12 163.32 244.12
FedProx 9.57 9.78 27.87 30.45 64.52 67.71 78.01 82.32 154.18 234.41
AFL 8.92 9.01 26.21 28.91 63.45 60.91 70.96 78.43 145.76 230.65
TiFL 8.21 8.47 25.43 26.22 50.24 53.08 61.23 67.12 128.94 190.82
Oort 7.84 7.52 20.13 23.01 41.89 48.65 53.78 56.45 108.38 168.23
Favor 9.51 9.44 25.62 27.32 58.98 59.71 65.78 71.07 149.45 195.64
FedMarl 9.16 8.82 21.60 23.42 38.23 42.54 51.08 54.09 119.83 150.76
FedGCS 6.13 6.48 15.34 18.76 32.28 40.19 43.98 44.42 94.59 138.71
Table 3: The average latency (min) of each training round for different selection approaches. FedGCS achieves the best performance across all the datasets.
  • 1
(a) MNIST (IID)
(b) MNIST (NonIID)
(c) CIFAR10 (IID)
(d) CIFAR10 (NonIID)
(e) CINIC10 (IID)
(f) CINIC10 (NonIID)
Figure 9: ToA: time to accuracy. Efficiency comparison of various schemes.
B.2 Baseline Details
We compare FedGCS with three types of representative device selection approaches, including
  • Random Selection:
    • FedAvg McMahan et al. (2017), a vanilla framework for FL without any operation. In each round $t$, the server randomly selects a subset of available devices $K_{t}$ from the total devices $K$. Each chosen device $k$ trains the model on its local dataset using the current global model parameters $w_{t}$, resulting in updated local parameters $w_{t+1}^{k}$. The server aggregates these updated local parameters using the formula:
      $w_{t+1}=\frac{1}{K_{t}}\sum_{k=1}^{K_{t}}w_{t+1}^{k}$ (14)
      This step effectively averages the local updates to form the new global model parameters.
    • FedProx Li et al. (2021) improves FedAvg by handling heterogeneous data and systems, adding a proximal term to the local training objective to address non-IID data and system differences. In each round $t$, a subset of devices $K_{t}$ is selected to train on their local data with the global model parameters $w_{t}$ and a proximal term. The local update $w_{t+1}^{k}$ is obtained by minimizing $L_{k}(w)+\frac{\mu}{2}\|w-w_{t}\|^{2}$, where $L_{k}(w)$ is the local loss, and $\mu$ adjusts the proximal term’s influence. FedProx’s main advantage is its ability to stabilize training across diverse data and system environments by ensuring local updates remain close to the global model.
  • Heuristic-based Selection:
    • AFL Goetz et al. (2019) uses a device selection method conditioned on the model and client data to enhance efficiency in each round $t$. This method focuses on devices likely to offer significant model improvements based on data diversity, model uncertainty, or past updates’ impact. Post-training, devices submit their updates and informativeness measures (like loss or uncertainty) to the server. The server then combines these updates, potentially weighted by informativeness, to revise the global model to $w_{t+1}$, the following:
      $w_{t+1}=\sum_{k=1}^{K_{t}}\alpha_{k}w_{t+1}^{k}$ (15)
      Here, $\alpha_{k}$ denotes the weight of the $k$-th device’s contribution.
    • TiFL Chai et al. (2020) groups devices by response latency to mitigate system heterogeneity, sorting them into tiers by capability, bandwidth, and data quality. Each tier has a Tier Server (TS) where selected devices contribute to local updates. These updates are first aggregated within tiers at the TS using:
      $w_{t+1}^{TS}=\frac{1}{K_{t}}\sum_{k=1}^{K_{t}}w_{t+1}^{k}$ (16)
      Then, Tier Servers forward their updates to a central server for global aggregation, potentially weighted by tier attributes:
      $w_{t+1}=\sum_{i=1}^{T}\beta_{i}w_{t+1}^{TS_{i}}$ (17)
      This balances individual contributions and overall data characteristics.
    • Oort Lai et al. (2021) optimizes device selection by integrating training loss and latency into a user-defined utility. To enhance efficiency, it’s crucial to maximize the per-unit-time statistical utility. The utility of client $i$ is defined as a product of her statistical utility and the global system utility, considering the duration of each training round, as shown in the equation:
      $\operatorname{Util}(i)=\underbrace{\left|B_{i}\right|\sqrt{\frac{1}{\left|B_{i}\right|}\sum_{k\in B_{i}}\operatorname{Loss}(k)^{2}}}_{\text{Statistical utility }U(i)}\times\underbrace{\left(\frac{T}{t_{i}}\right)^{\textbf{1}\left(T<t_{i}\right)\times\alpha}}_{\text{Global sys utility }}$ (18)
    Dataset &Model CV Tasks NLP Tasks
    MNIST CIFAR10 CINIC10 TinyImageNet Shakespeare Wikitext
    LeNet5 Resnet18 VGG16 ShuffleNet LSTM GPT-2
    FedAvg 0.025 0.021 3.42 3.02 4.65 4.42 8.92 9.14 16.28 23.67
    FedProx 0.023 0.020 3.13 3.06 4.43 4.39 8.87 9.02 15.89 23.45
    AFL 0.019 0.019 2.69 2.66 4.02 4.11 8.31 8.54 14.12 21.89
    TiFL 0.018 0.016 2.31 2.01 3.56 3.31 6.32 6.73 13.25 20.56
    Oort 0.014 0.015 1.91 1.86 3.23 3.02 6.80 6.54 13.17 18.32
    Favor 0.021 0.019 2.79 2.81 4.51 4.28 8.62 8.55 15.96 23.02
    FedMarl 0.012 0.014 1.82 1.82 3.17 2.98 6.72 6.29 12.56 19.26
    FedGCS 0.007 0.010 1.71 1.51 2.26 2.71 5.90 6.02 11.82 17.85
    Table 4: The average energy consumption (KJ) of each training round for different selection approaches. FedGCS achieves the best performance across all the datasets.
    • 1
    (a) MNIST (IID)
    (b) MNIST (NonIID)
    (c) CIFAR10 (IID)
    (d) CIFAR10 (NonIID)
    (e) CINIC10 (IID)
    (f) CINIC10 (NonIID)
    Figure 10: ToE: time to energy. Efficiency comparison of various schemes.
  • Learning-based Selection:
    • Favor Wang et al. (2020) utilizes accuracy to determine local weight selection, mitigating non-IID impacts and enhancing convergence. It involves training a DRL agent via a double-deep Q-learning Network (DQN). Despite evident disparities in local model weights, which hold guiding data for device selection, the DQN agent’s training aims to optimize the expected total discounted reward, represented as $R=\sum_{t=1}^{T}\gamma^{t-1}r_{t}=\sum_{t=1}^{T}\gamma^{t-1}\left(\Xi^{\left(\omega_{t}-\Omega\right)}-1\right)$.
    • FedMarl Zhang et al. (2022) applies multi-agent reinforcement learning to select devices for federated learning. It strategically selects RL agents each round to optimize accuracy, training latency, and communication cost. The reward $r_{t}$ for each round $t$ is given by:
      $r_{t}=w_{1}[U(Acc(t))-U(Acc(t-1))]-w_{2}H_{t}-w_{3}B_{t}$ (19)
      where $H_{t}$, the processing latency, is:
      $H_{t}=\max_{1\leq n\leq N}\left(H_{t,n}^{p}\right)+\max_{n:1\leq n\leq N,a_{n}^{t}=1}\left(H_{t,n}^{rest}+H_{t,n}^{u}\right)$ (20)
      In this, $\max_{1\leq n\leq N}H_{t,n}^{p}$ signifies the maximum time for generating probing losses, used by MARL agents for client selection and model update. The time for client $n$ to complete training and upload updates is $\max_{n:1\leq n\leq N,a_{n}^{t}=1}\left(H_{t,n}^{\text{rest }}+\right$. $\left.H_{t,n}^{u}\right)$ Here, $U(.)$ is a utility function ensuring even modest improvements in $Acc(t)$ are recognized towards the end of the learning process, and $B_{t}$ represents the total communication cost.
B.3 Implementation Details
We ran Oor, Favor, and FedMarl 100 times each (for a total of 300 runs) to collect “selection-score” pair data to be used as training data for the subsequent model. For the data augmentation part, we randomly shuffled each selection 25 times to model the order-independent properties of sets. We adopt a single-layer LSTM as the Encoder and Decoder backbones and use two-layer feed-forward networks to implement the Evaluator. The hidden state sizes for the Encoder, Decoder, and Evaluator are 64, 64, and 200, respectively. The embedding size of each device ID token is set to 32. To train FedGCS, we set the batch size as 1024, the learning rate as 0.001, and $\alpha$ to 0.8. We utilize the top-25 device selection records as starting points for gradient-based optimization. For FL part, we set the number of devices that are available in the device pool as $J=100$, while $T=10$ devices are selected to participate in each round and $T=20$ for FedMarl for further action-making. The number of training rounds $r$ is set to $r=10$ for GPT-2, and $r=50$ for others. The clients perform local training for $ep=5$ epochs in each training round.
Appendix C Additional Results
C.1 Overall Performance of Latency
Table 3 shows the total processing latency for all the methods. Figure 9 shows the results of Time to Accuracy (ToA) in MNIST (IID), MNIST (Non-IID), CIFAR10 (IID), CIFAR10 (Non-IID), CINIC10 (IID), CINIC10 (Non-IID).
C.2 Overall Performance of Energy Consumption
Table 4 shows the total energy cost for all the methods. Figure 10 shows the results of Energy to Accuracy (EoA) in MNIST (IID), MNIST (Non-IID), CIFAR10 (IID), CIFAR10 (Non-IID), CINIC10 (IID), CINIC10 (Non-IID).